# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@ 4.2. Local-energy-preserving approximation
Observe that (4.9) has better regularity in nonlinear terms, and one could obtain the existence of solutions via a similar argument using the Galerkin method and Banach fixed point theorem as in [42]. We remark that this kind of approximation by regularization was introduced by Leray [32] for the Navier–Stokes equations. In [7], Caffarelli, Kohn, and Nirenberg adopt a similar approximation through a “retarded mollification” to obtain the existence of suitable weak solutions for Navier–Stokes equations. It is also applicable to more complicated hydrodynamic equations (see for instance, [13, 15]). The key feature of this type of approximation is that it preserves global and local energy inequalities.
Multiplying (4.9)${}_{1}$ by ${{\bf{v}}}^{\sigma}$, and (4.9)${}_{3}$ by $-\Delta{{\bf{d}}}^{\sigma}+f({{\bf{d}}}^{\sigma})$, integrating over $\mathbb{T}^{3}$ and adding two equations together we obtain (see also Corollary 4.2)
@@
(4.10) $\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}}x+\int_{\mathbb{T}^{3}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}){\,\mathrm{d}}x$
$=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}+{{\bf{z}}}\cdot\nabla\Phi[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})){\,\mathrm{d}x}.$
By the Hölder inequality and Young inequality we can estimate
$\displaystyle\left|\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot\nabla\Phi[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})){\,\mathrm{d}x}\right|$
$\displaystyle\leq\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2},$
$\displaystyle\leq\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}$
$\displaystyle\leq\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2},$
where $C$ is a constant which is independent of $\sigma$, and
@@
$\displaystyle\leq\int_{\mathbb{T}^{3}}|{{\bf{z}}}||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|+|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|$
$\displaystyle\leq\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}+\|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}$
@@
$\displaystyle\leq\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2}\right)$
@@
$\displaystyle\leq\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}\right)$
$\displaystyle\leq\frac{1}{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}\right).$
Meanwhile, by a direct calculation, we can show that

### v2

@@ 4.1 Energy-inequality-preserving approximation
As the key feature of this type of approximations is that global and local energy inequalities are preserved, it is also applicable to more complicated hydrodynamic equations (see, for instance, [19, 21]). Moreover, we would like to point out that, for the Ericksen stress tensor term, we have employed the same observation as discussed in Remark 1.
We note that (17) has better regularity in nonlinear terms, and one could obtain the existence of solutions globally strong in the PDE sense via a similar argument using the Galerkin method and Banach fixed point theorem as in [26] and [79]. More precisely, for each fixed $\sigma$, we have
$\displaystyle\begin{cases}&{{\bf{v}}}^{\sigma}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3}))\cap L^{\infty}(0,T;V),\quad\partial_{t}{{\bf{v}}}^{\sigma}\in L^{2}(0,T;H),\\ &{{{\bf{d}}}}^{\sigma}\in L^{2}(0,T;H^{2}(\mathbb{T}^{3}))\cap L^{\infty}(0,T;V),\quad\partial_{t}{{{\bf{d}}}}^{\sigma}\in L^{2}(0,T;H),\\ &{\pi}^{\sigma}\in L^{2}(0,T;H^{1}(\mathbb{T}^{3})).\end{cases}$ (18)
Moreover, we note that subtracting any constant from $\pi^{\sigma}$ yields again a solution to (17). So by subtracting ${\pi}^{\sigma}$ with its mean, $\int_{\mathbb{T}^{3}}{\pi}^{\sigma}$, we will obtain a function with zero mean; we will choose this function as the solution. Thus, without loss of generality, from now on we will assume that
$\int_{\mathbb{T}^{3}}{\pi}^{\sigma}=0.$ (19)
4.1.1 Pathwise estimates independent of the approximation parameter
We can derive the following estimates on $({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma})$ independent of $\sigma$, with a fixed arbitrary $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies the bound specified in (13); these estimates will be used to achieve the compactness argument later on (see Section 4.2.1).
Lemma 3.
[Energy estimates independent of $\sigma$] Fix $\omega\in\Omega$ such that ${{\bf{z}}}(\omega)\in L_{\rm loc}^{\infty}(\mathbb{T}^{3}\times[0,\infty))$. Consider the triple $({{\bf{v}}}^{\sigma}(\omega),{\pi}^{\sigma}(\omega),{{\bf{d}}}^{\sigma}(\omega))$, a classical strong (in the PDE sense) solution to (17), then we have the following estimates:
@@ 4.1.1 Pathwise estimates independent of the approximation parameter
$\begin{cases}&{\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;H)}^{2}+\|{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;H^{1}(\mathbb{T}^{3}))}^{2}+\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}}\\ &\leq\Psi(T),\\ &\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C\Psi(T)+{C\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}},\\ &\|\partial_{t}{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}\leq C(\sqrt{\Psi(T)}+\Psi(T)),\\ &\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C(\Psi(T)+T^{\frac{1}{10}}\sqrt{\Psi(T)}),\\ &|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}_{L^{2}(Q_{T})}\leq C\Psi(T),\end{cases}$ (20)
where $C$ is a constant that is independent of $\sigma$, and
$\Psi(T)=$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\|{{\bf{z}}}(t)\|_{L^{4}(Q_{T})}^{4}\right)e^{C\int_{0}^{T}{\frac{1}{2}}+\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}.$ (21)
Proof.
The proof of this lemma is based on deriving suitable energy estimates. Due to the length of the derivation of the global energy equality, we have placed it in Appendix B. In particular, all of the estimates stated above will be derived from equation (86) in Corollary 14.
Firstly, we deal with the two terms on the right-hand-side (RHS) of (86) one by one. For the first term we have the following estimation:
@@
$\displaystyle\leq$ $\displaystyle\int_{\mathbb{T}^{3}}|{{\bf{z}}}||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|+|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]||\nabla{{\bf{v}}}^{\sigma}||{{\bf{z}}}|$
$\displaystyle\leq$ $\displaystyle\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}+\|\Phi_{\sigma}[{{\bf{v}}}^{\sigma}]\|_{L^{4}(\mathbb{T}^{3})}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}$
$\displaystyle\leq$ $\displaystyle\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2}\right)$
$\displaystyle\leq$ $\displaystyle\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{{H^{1}(\mathbb{T}^{3})}}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}\right)$
$\displaystyle\leq$ $\displaystyle\frac{1}{4}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}+{\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)$
@@
$\displaystyle\leq$ $\displaystyle\frac{1}{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+{\left(\dfrac{1}{2}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+\dfrac{1}{2}\right)\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)$
@@
$\displaystyle\leq$ $\displaystyle\frac{1}{2}\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{3}{2}}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right),$
where $C$ is a constant [^1] Note that C may be different at each occurrence below but all of them are independent of $\sigma$. which is independent of $\sigma$. For the second term, by the Hölder inequality and Young inequality, we obtain the following estimation: